Caffeine:following are the number of grams of carbohydrates in 12 oz expression beverages offered at coffee shop

45 29 13 61 18 33 21 40 41 26 27 12

30 14 19 34 22 25 42 11 47 57 32 32

Find the first and third quartile of the data.

Find the mean of the data.

Find the upper and Lower outliner boundaries.
Answer #1

putting above data in ascending order:

x
11
12
13
14
18
19
21
22
25
26
27
29
30
32
32
33
34
40
41
42
45
47
57
61

first quartile =Q1= npth value =24*0.25 th value =6 th value =(19+21)/2 =20

third   quartile =Q3= npth value =24*0.75 th value =18 th value =(40+41)/2 =40.5

mean of the data =sum of data values/n=731/24 =30.46

upper oulier boundary =Q3+1.5*(Q3-Q1)=40.5+1.5*(40.5-20)=71.25

lower oulier boundary =Q1-1.5*(Q3-Q1)=20-1.5*(40.5-20)=-10.75